Point charges q1 = 50 µC and q2 = −75 µC are placed 1 m apart. Determine the

magnitude and direction of the force on a third charge q3 = 20 µC placed midway

between q1 and q2


1
Expert's answer
2021-09-28T11:11:47-0400

The force from q1:


"F_{13}=k\frac{q_1q_3}{r^2},\\\space\\
F_{23}=k\frac{q_2q_3}{r^2},\\\space\\
F_3=F_{23}+F_{13}=\frac k{r^2}(q_2q_3+q_1q_3)=\\\space\\
=\frac {9·10^9}{0.5^2}((-75\mu)(20\mu)+(50\mu)(20\mu))=\\\space\\
=-18\text{ N toward q}_1\text{ or 18 N toward q}_2."

The third charge will be pulled toward q2 (pushed from q1 toward q2).